OpenStudy (anonymous):

Multiply the rational expressions. (15/r^(2)-4r+4) X (r^(2)-4/30r^(2))

OpenStudy (anonymous):

\[\frac{ 15 }{ r ^{2}-4r+4 } \times \frac{ r ^{2}-4 }{ 30r ^{2} }\]

OpenStudy (e.mccormick):

OK. Where did this giive you problems?

OpenStudy (e.mccormick):

And one early thing to do with this one is to factor both sides and d some canceling. That will make the multiplication easier.

OpenStudy (anonymous):

I just want to be sure if its right. I got r+2/r^2(2r-4)

OpenStudy (e.mccormick):

AH. OK. Let me finish it then...

OpenStudy (e.mccormick):

If you are going to distribute the 2 on the bottom, you would also distribute the \(r^2\). As for the numbers you go, they are correct.

OpenStudy (anonymous):

Okay

OpenStudy (e.mccormick):

And I do not see anything else that would cancel, so you are in simplified form.

OpenStudy (anonymous):

I see. So what if I got r^2-4/2r^2(r^2-4r+4) ?

OpenStudy (e.mccormick):

That looks like it would have been without the canceling, which is not reduced form. They generally want things as simplified as possible. So on a hand grading, it would be partial credit at best.